OpenStudy (anonymous):

Simplify. (2x+3)(3x-5)

OpenStudy (rosedewittbukater):

6x^2 -10x +9x - 15 6x^2 - x - 15

OpenStudy (kainui):

Did you learn FOIL? It means multiply the First terms, Outer, Inner, and Last terms and add them all up.

OpenStudy (anonymous):

(2x+3)(3x-5) =>2x*(3x-5) + 3*(3x-5) 6x^2 - 10x + 9x - 15 6x^2 - x - 15

OpenStudy (anonymous):

explain more pls

hero (hero):

Do you know distributive rule?

hero (hero):

a(b + c) = ab + ac

hero (hero):

Well if we let: a = 2x+3 b = 3x c = -5 Then: (2x + 3)(3x - 5) becomes (2x + 3)(3x) +(2x + 3)(-5)